i really need youre help with this problem
how many ounces of a 50% acetic solution should a photographer add to 32 ounces of a 5% acetic acid solution to obtain a 10% acetic acid solution?

Answer by stanbon(75887)

how many ounces of a 50% acetic solution should a photographer add to 32 ounces of a 5% acetic acid solution to obtain a 10% acetic acid solution?
-------------------
Equation:
acid + acid = acid
-------
0.50x + 0.05*32 = 0.10*(32+x)
Multiply thru by 100 to get:
50x + 5*32 = 10*32 + 10x
40x = 5*32
x = 4 ounces (amt. of 50% solution needed in the mixture)
======================================
Cheers,
Stan H.